Let F be the R²-R function defined by f(x,y)=Inxy and let r be the R-R² function defined by r(t)=(e^t;t).


1.determine the composite function F o r: (simplify your answer).


2.determine gradf (x,y) and r'(t).


3.determine the derivative function (f o r)' by


3.1.differentiating the expression obtained in (1).


3.2.using the chain rule (theorem ) compare your answer.


1
Expert's answer
2022-04-19T04:47:48-0400

"1:\\\\f\\circ r\\left( x \\right) =\\ln \\left( e^t\\cdot t \\right) \\\\2:\\\\gradf\\left( x,y \\right) =\\left[ \\begin{array}{c}\tf'_x\\\\\tf'_y\\\\\\end{array} \\right] =\\left[ \\begin{array}{c}\t\\frac{1}{x}\\\\\t\\frac{1}{y}\\\\\\end{array} \\right] \\\\r'\\left( t \\right) =\\left[ \\begin{array}{c}\t\\left( e^t \\right) '\\\\\tt'\\\\\\end{array} \\right] =\\left[ \\begin{array}{c}\te^t\\\\\t1\\\\\\end{array} \\right] \\\\3.1:\\\\\\left( f\\circ r \\right) '\\left( t \\right) =\\left( t+\\ln t \\right) '=1+\\frac{1}{t}\\\\3.2:\\\\\\left( f\\circ r \\right) '\\left( t \\right) =\\frac{\\partial f}{\\partial x}\\left( r_1\\left( t \\right) ,r_2\\left( t \\right) \\right) \\frac{\\partial r_1}{\\partial t}+\\frac{\\partial f}{\\partial y}\\left( r_1\\left( t \\right) ,r_2\\left( t \\right) \\right) \\frac{\\partial r_2}{\\partial t}=\\\\=\\frac{1}{r_1\\left( t \\right)}e^t+\\frac{1}{r_2\\left( t \\right)}\\cdot 1=\\frac{e^t}{e^t}+\\frac{1}{t}=1+\\frac{1}{t}\\\\Answers\\,\\,are\\,\\,same"